# Adaptive Reduced Multilevel Splitting

### v1

@@ 6 Conclusions and Perspectives
As $S^{(k)}$ gains in accuracy and $l^{(k)}$ increases, the reduced score approximation should be refined using samples of the state space in regions closer and closer to the rare event of interest. However, an essential ingredient in achieving this desirable convergence is to determine a heuristic for setting the first non-achievable level $l^{(k)}$ suitable for the accuracy of the current reduced score approximation $S^{(k)}$. In addition, one of the conditions of importance sampling is to ensure that the proposal dominates the target. Justified by theoretical arguments, these two problems are addressed in ARMS using entropy constraints: $l^{(k)}$ is chosen such that $i)$ the relative entropy between a pessimistic target and the proposal corresponds to a given logarithmic cost of importance sampling; $ii)$ the relative entropy between an optimistic target and the proposal remains finite. Pessimistic and optimistic targets are constructed using a quantification of the worst-case error associated with the reduced score approximation.
A final point addressed by ARMS is the adaptation of the proposal to a new approximation of the reduced score $S^{(k+1)}$. In particular, the algorithm ensures that the updated proposal dominates a well-chosen previous proposal, so that the two proposals can be bridged in a single AMS simulation step. To achieve this, ARMS searches among past simulation indexes $k_{bridge}<k+1$ for an approximation $S^{(k_{bridge})}$ and an initial level $l_{bridge}(k+1)$ such that $iii)$ the relative entropy between the new proposal $\mu_{l_{\mathrm{bridge}}(k+1)}^{(k+1)}$ at its initial level and the past proposal ${\mu_{l^{(k_{\mathrm{bridge}})}}^{(k_{\mathrm{bridge}})}}$ at its final level is kept below a certain standard AMS threshold and, in addition, such that both entropy constraints $i)$ and $ii)$ are satisfied.
@@
In addition to demonstrating the unbiasedness of our estimator in an idealized setting, we evaluate in our numerical experiments the empirical convergence of the ARMS algorithm on a toy model and a rare-event problem based on a more realistic PDE. Our study confirms that entropy criteria play a crucial role in ARMS. In particular, as expected, too high a value of the logarithm of the cost in the entropy criterion $i)$ can take ARMS into spurious regions of the state-space remote from the region of interest, resulting in a large variance of the IS estimator. Too small a value gives a more desirable variance but a slower rate of convergence. We show empirically that the approximation procedure using the entropy criterion $iii)$ significantly reduces simulation cost while achieving a comparable squared error.
@@
The prospects arising from this work are numerous. These include answering open theoretical questions, including the consistency of ARMS, which is proven in the present paper only in a formal sense. No straightforward answers seem available, as we are confronted with a non-standard framework in which interacting particles in AMS simulations have distributions with varying supports which depend on random approximations of the reduced score. Future research directions also include extending ARMS to smooth proposal distributions, with the advantage of removing nesting and domination constraints, or particularizing this approach to the simulation of rare events driven by stochastic dynamics.
@@ A.1 Generic quantification rules
Unfortunately, importance sampling becomes quickly infeasible if the target distribution diverges too widely from the proposal, a typical phenomenon in high dimension. This can be readily seen by computing the relative variance of the importance sampling estimator for $\varphi=\,{\rm l}{\rm l}$:
$\operatorname{\mathbb{V}ar}_{\mu}(f(X))/p^{2}=\eta(\frac{\mathrm{d}\eta}{\mathrm{d}\mu})-1=\mathrm{e}^{\mathrm{Ent}_{2}(\eta\mid\mu)}-1,$
@@
As is thoroughly argued in [7] (see also [6]), the variance ($e^{\mathrm{Ent}_{2}(\eta\mid\mu)}-1$) is in general a pessimistic quantification of importance sampling, whose cost is better estimated using a $\mathbb{L}^{1}$-error, or a success probability within a tolerance precision. For a target $\eta$ with proposal $\mu$, it has been proven in  [7] that the cost of importance sampling – in terms of sample size $K$– is roughly given (up to a reasonable tail condition on the log-density $\ln\frac{\mathrm{d}\eta}{\mathrm{d}\mu}$), by the exponential of the relative entropy between $\eta$ and $\mu$:
$K\simeq\mathrm{e}^{\mathrm{Ent}(\eta\mid\mu)}.$
@@
shows that indeed, variance is (in general) not a sharp estimate of the required sample size for importance sampling.
@@ Appendix B Consistency of the Estimator
For simplicity (but the general case is similar), we restrict to the case where $\tau^{(k)}=0$ for $k\geq K_{\mathrm{min}}+1$ in the estimator (3.1), as well as $\varphi=\,{\rm l}{\rm l}$. We thus consider the rare event probability estimator: $\hat{p}_{l}^{H}=\hat{\gamma}^{H}_{l}(\,{\rm l}{\rm l}).$
@@
$\operatorname{\mathbb{V}ar}\left[\hat{p}_{l}^{H}\right]/{p^{\ast}_{l}}^{2}=\frac{1}{H^{2}}\sum_{k=K_{\mathrm{min}}+1}^{K_{\mathrm{min}}+H}\left(\frac{\pi(\,{\rm l}{\rm l}_{S^{(k)}>l^{(k)}})}{p^{\ast}_{l}}-1\right).$ (B.1)
@@
We first show that the random sequence
$H\mapsto M_{H}\displaystyle\mathop{=}^{{\rm def}}H\,(\hat{\gamma}_{l}^{H}-\gamma^{\ast}_{l}),$
is a martingale with respect to the filtration generated by the snapshots. Indeed, it follows from the definition (3.1) that
$\mathbb{E}[M_{H}|X_{1},\ldots,X_{H}]$ $=\mathbb{E}[\sum_{k^{\prime}=K_{\mathrm{min}}+1}^{K_{\mathrm{min}}+H}f_{l}^{(k^{\prime})}(X_{k^{\prime}+1})\delta_{X_{k^{\prime}+1}}-h\gamma^{\ast}_{l}|X_{1},\ldots,X_{H}],$
$=M_{H-1}+\mathbb{E}[f_{l}^{(\tilde{k})}(X_{\tilde{k}+1})\delta_{X_{\tilde{k}+1}}-\gamma^{\ast}_{l}|X_{1},\ldots,X_{H}],$
$=M_{H-1}+\pi(\,{\rm l}{\rm l}_{S^{\ast}>l})-\gamma^{\ast}_{l}=M_{H-1},$
where $\tilde{k}={K_{\mathrm{min}}+H}$.
Now, since $M_{H}$ is a martingale, the estimators are unbiased $\mathbb{E}\,\hat{\gamma}_{l}^{H}=\gamma_{l}^{\ast},$ and their variances are given by the average of their quadratic variation. In particular, the variance is for the rare event probability estimator
$H^{2}\operatorname{\mathbb{V}ar}\left[\hat{p}_{l}^{H}\right]$ $={p^{\ast}_{l}}^{2}\sum_{k=K_{\mathrm{min}}+1}^{K_{\mathrm{min}}+H}\left(\mathrm{e}^{\mathrm{Ent}_{2}\left(\eta_{l}^{\ast}\mid\mu^{(k)}_{l^{(k)}}\right)}-1\right),$

### v3

@@ 5 Conclusions and Perspectives
As $S^{(k)}$ gains in accuracy and $l^{(k)}$ increases, the reduced score approximation should be refined using samples of the state space in regions closer and closer to the rare event of interest. However, an essential ingredient in achieving this desirable convergence is to determine a criteria for setting the first non-achievable (in the sense of Importance Sampling cost) level $l^{(k)}$ by the accuracy of the current reduced score approximation $S^{(k)}$. In addition, one must ensure that the support of the proposal always includes the final rare event target. Justified by theoretical arguments, these two problems are addressed in ARMS using entropy constraints: $l^{(k)}$ is chosen such that $i)$ the relative entropy between a pessimistic rare event target at level $l^{(k)}$ and the proposal corresponds to a given logarithmic cost of importance sampling; $ii)$ the relative entropy between an optimistic final rare event target and the proposal remains finite. Pessimistic and optimistic rare event targets are constructed using a quantification of the worst-case error associated with the reduced score approximation.
A final point addressed by ARMS is the sampling of the proposal given by a new approximation of the reduced score $S^{(k+1)}$. In particular, the algorithm finds a (usually lower) level $l_{b}(k+1)$ at which the updated proposal is dominated (inclusion in terms of support) by a well-chosen previous (carefully recorded) proposal, so that the updated proposal can be sampled by ’bridging’ quickly from the recorded proposal in a single AMS simulation step. In addition, the updated proposal at level $l_{b}(k+1)$ must satisfy entropy constraints $i)$ and $ii)$ of the previous paragraph.
@@
In addition to demonstrating the unbiasedness of our estimator in an idealized setting, we evaluate in our numerical experiments the empirical convergence of the ARMS algorithm on a toy model and a rare-event problem based on a more realistic PDE. Our study confirms that entropy criteria play a crucial role in ARMS. In particular, as expected, too high a value of the logarithm of the cost in the entropy criterion $i)$ can take ARMS into spurious regions of the state-space remote from the region of interest, resulting in a large variance of the IS estimator. Too small a value gives a more desirable variance but a slower rate of convergence. We show empirically that the approximation procedure using the entropy criterion $iii)$ significantly reduces simulation cost (around a factor $10$) while achieving a comparable squared error. In the same way in both experiments, we observe a squared error gain of the order of a decade for a given simulation cost. Our numerical experiments show that bridging is a good option for small computational budgets, with a saving of around 2 compared with restarting the simulation at each iteration. However, for larger computational budgets, both procedures appear to be equivalent in terms of cost multiplied by squared error.
@@
The prospects arising from this work are numerous. These include answering open theoretical questions, including the consistency of ARMS, which is proven in the present paper only in a formal sense. No straightforward answers seem available, as we are confronted with a non-standard framework in which interacting particles in AMS simulations have distributions with varying supports which depend on random approximations of the reduced score. Future research directions also include extending ARMS to smooth proposal distributions, with the advantage of removing nesting and domination constraints and the possibility of being more flexible in the choice of error estimates that need not have strict error bounds.
@@ A.1 Generic quantification rules
Unfortunately, importance sampling becomes quickly useless if the target distribution diverges too widely from the proposal, a typical phenomenon in high dimension. This can be readily seen by computing the relative variance of the importance sampling estimator for $\varphi=\,{\rm l}{\rm l}$:
$\operatorname{\mathbb{V}ar}_{\mu}(f(X))/p^{2}=\eta(\frac{\mathrm{d}\eta}{\mathrm{d}\mu})-1=\mathrm{e}^{\mathrm{Ent}_{2}(\eta\mid\mu)}-1,$ (A.1)
@@
As is thoroughly argued in [8] (see also [7]), the variance ($e^{\mathrm{Ent}_{2}(\eta\mid\mu)}-1$) is in general a pessimistic quantification of importance sampling, whose cost is better estimated using a $\mathbb{L}^{1}$-error, or a success probability within a tolerance precision. For a target $\eta$ with proposal $\mu$, it has been proven in  [8] that the cost of importance sampling – in terms of sample size $K$– is roughly given (up to a reasonable tail condition on the log-density $\ln\frac{\mathrm{d}\eta}{\mathrm{d}\mu}$), by the exponential of the relative entropy between $\eta$ and $\mu$: $K\simeq\mathrm{e}^{\mathrm{Ent}(\eta\mid\mu)}.$
@@
shows that indeed, variance is a more pessimistic estimator of the required sample size for importance sampling.
@@ Appendix B Consistency of the Estimator
This section analyses the idealized algorithm obtained with $N=+\infty$ many particles (clones) in ARMS;
For simplicity (but the general case is similar), we study the estimator (2.9) in the case of a test function $\varphi=\,{\rm l}{\rm l}$. We thus consider the rare event probability estimator: $\hat{p}_{l}^{H}=\hat{\gamma}^{H}_{l}(\,{\rm l}{\rm l}).$
@@
$\operatorname{\mathbb{V}ar}\left[\hat{p}_{l}^{H}\right]/{p^{\ast}_{l}}^{2}=\frac{1}{H^{2}}\sum_{k=K_{j_{0}}+1}^{K_{j_{0}}+H}\left(\frac{\mathbb{E}({Z_{l^{(k)}}^{(k)}})}{p^{\ast}_{l}}-1\right),$ (B.1)
where the random variable ${Z_{l^{(k)}}^{(k)}}$ is a function of the past snapshots.
@@
We first show that for any test function $\varphi$ the random sequence $H\mapsto M^{\varphi}_{H}\displaystyle\mathop{=}^{{\rm def}}H\,(\hat{\gamma}_{l}^{H}(\varphi)-\gamma^{\ast}_{l}(\varphi)),$ is a martingale with respect to the filtration generated by the snapshots. Indeed, defining this filtration by $\mathcal{F}_{H}=\sigma(X_{1},\ldots,X_{K_{j_{0}}+1+H})$, it follows from the definition (2.9) that
$\mathbb{E}[M^{\varphi}_{H}|\mathcal{F}_{H-1}]$ $=\mathbb{E}[\sum_{k=K_{j_{0}}+1}^{K_{j_{0}}+H}f_{l}^{(k)}(X_{k+1})\varphi({X_{k+1}})-H\gamma^{\ast}_{l}(\varphi)|\mathcal{F}_{H-1}],$
$=M^{\varphi}_{H-1}+\mathbb{E}[f_{l}^{({K_{j_{0}}+H})}(X_{{K_{j_{0}}+H}+1})\varphi({X_{{K_{j_{0}}+H}+1}})-\gamma^{\ast}_{l}(\varphi)|\mathcal{F}_{H-1}],$
$=M^{\varphi}_{H-1}+\pi(\varphi\,{\rm l}{\rm l}_{S^{\ast}>l})-\gamma^{\ast}_{l}(\varphi)=M^{\varphi}_{H-1}.$
Now, since $M^{\varphi}_{H}$ is a martingale, the estimator is unbiased $\mathbb{E}\,\hat{\gamma}_{l}^{H}(\varphi)=\gamma_{l}^{\ast}(\varphi),$ and its variance is given by its quadratic variations given by the sum over $k$ of the local variances. In particular, using (A.1), the variance of the rare event probability estimator is then
$\operatorname{\mathbb{V}ar}\left[\hat{p}_{l}^{H}\right]/{p^{\ast}_{l}}^{2}$ $=\frac{1}{H^{2}}\sum_{k=K_{j_{0}}+1}^{K_{j_{0}}+H}\left(\mathbb{E}(\mathrm{e}^{\mathrm{Ent}_{2}\left(\eta_{l}^{\ast}\mid\mu^{(k)}_{l^{(k)}}\right)})-1\right),$